Calculate the number of gallons of gasoline that one would have to obtain the same amount of energy that could be recovered from the conversion of the entire mass of one drop to energy by nucleus. Gasoline density = 745 kg/m3 and it yeilds 18900 Btu/lb when burned.
1
Expert's answer
2020-09-28T08:06:21-0400
"E=mc^2=\rho V c^2=(745)(61.6\cdot10^{-9})(3\cdot10^8)^2=4.13\cdot10^{12}J"

"E=QM=Q\rho V\\4.13\cdot10^{12}=18900\cdot 2326\cdot 745\cdot V"

"V=126\ m^3=33286\ gal"